OpenStudy (lochana):

\((x+y)(dx-dy) = dx + dy\) what to do when equations are parallel to each other.

OpenStudy (fibonaccichick666):

Have you tried distributing then seeing if exact?

OpenStudy (lochana):

yes. I tried exact method.but didn't work for me

OpenStudy (lochana):

after re arranging it. I get\[\frac{dy}{dx} = \frac{x+y-1}{x+y+1}\]

OpenStudy (fibonaccichick666):

\[(x+y)(dx-dy) = dx + dy\]\[xdx+ydx -xdy-ydy = dx + dy\] \[(x+y-1)dx -(x+y+1) dy=0

OpenStudy (fibonaccichick666):

\[(x+y-1)dx -(x+y+1) dy=0\]**

OpenStudy (lochana):

yeah. exact doesn't work. right?

OpenStudy (fibonaccichick666):

It is one of these http://mathworld.wolfram.com/dAlembertsEquation.html

ganeshie8 (ganeshie8):

if it is not exact, you may try making it exact

OpenStudy (lochana):

and x+y-1 and x+y+1 are parallel. so solutions doesn't exist too

OpenStudy (lochana):

@ganeshie8 what do yo mean?

OpenStudy (fibonaccichick666):

Did you check out the link? You should be able to get it from there. We can find another way if you haven't learned that yet.

OpenStudy (lochana):

give me a sec. I will read it

OpenStudy (lochana):

seems confusing. can you explain it.

OpenStudy (freckles):

make it separable with the sub v=x+y

OpenStudy (lochana):

@freckles seems working. wait a minute

OpenStudy (fibonaccichick666):

go with freckles, then you don't need PDE

OpenStudy (lochana):

@freckles thanks. It worked

OpenStudy (lochana):

@FibonacciChick666 thanks for helping

OpenStudy (fibonaccichick666):

sorry I wasn't better help. I always forget about substitution

OpenStudy (lochana):

\[v = y + x \\ y = v -x \\ \frac{dy}{dx} = \frac{dv}{dx} - 1 \ \rightarrow 1\]

OpenStudy (lochana):

\( \frac{dy}{dx} = \frac{x+v-x-1}{x+v-x+1} = \frac{v-1}{v+1}\)

OpenStudy (lochana):

\[\frac{dv}{dx} = \frac{2v}{v+1} \\ \int \frac{v+1}{2v}dv = \int dx\]

OpenStudy (lochana):

\[\frac{1}{2}v + \frac{1}{2}lnv = x + C \\ \frac{(x+y) + ln(y+x)}{2} = x + C\]

OpenStudy (lochana):

is that right?

OpenStudy (fibonaccichick666):

I'd solve this one for C, but that is what I got

OpenStudy (freckles):

looks good

OpenStudy (lochana):

Thanks guys. I was struggling on this more than 1 hour. haha..

OpenStudy (freckles):

i didn't know a sub would work it just look like it might once you put it in that one form you did the y'=(x+y-1)/(x+y+1) form

OpenStudy (freckles):

that's the cool thing about math you don't know what will work and sometimes you have to just try things

OpenStudy (lochana):

no. @ganeshie8 also said about it before. but he didn't continue sub method. It perfectly worked..

OpenStudy (lochana):

exactly